ANALYSIS OF ERROR PROPAGATION IN THE ABS CLASS FOR LINEAR SYSTEMS

A. GALANTAI

Department of Computer Science, The University of Agriculture, Godollo, 2103, Hungary

(Received May 15, 1989; revised March 12, 1990)

Abstract. Broyden’s backward error analysis technique is applied to evaluate the numerical stability of the ABS class of methods for solving linear systems.

Key words and phrases: Backward error analysis, optimally stable method.